# thanks: (No relevant financial or non-financial competing interests. Derived from preprint [23].)

### v3

@@ Abstract
The multi-armed bandit (MAB) problem is a widely studied model in the field of operations research for sequential decision making and reinforcement learning. This paper mainly considers the classical MAB model with the heavy-tailed reward distributions. We introduce the extended robust UCB policy, which is an extension of the pioneering UCB policies proposed by Bubeck et al. [5] and Lattimore [21]. The previous UCB policies require the knowledge of an upper bound on specific moments of reward distributions or a particular moment to exist, which can be hard to acquire or guarantee in practical scenarios. Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p$ and $q$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of the UCB policies for the heavy-tailed reward distributions.
keywords:
frequentist multi-armed bandit problems; the light-tailed reward distributions; the heavy-tailed reward distributions; upper confidence bounds
@@ 1 Introduction
The multi-armed bandit problem is an important topic in operations research where a player aims to identify and select the best arm to operate such that the long-term expected reward is maximized. The player’s action at each time involves balancing the trade-off between exploration and exploitation, i.e., choosing an arm to learn its statistics vs. choosing an arm appearing to be best based on past observations. Under resource constraints, the way we acquire knowledge becomes crucial for the efficiency in maximizing the long-term reward in modern AI systems [31]. While knowledge can be generally treated as data (usually not free in a strict sense), the struggle between exploration (sampling) and exploitation (consuming) as the prototypical dilemma in reinforcement learning is expressly reflected in an MAB model [28, 20, 3, 24, 5]. Specifically, the original stochastic model of the multi-armed bandit problem considers a bandit machine with several arms; each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or, agent). Each reward distribution has a mean, say $\theta_{i}$. The player who has limited knowledge about the underlying reward distributions (no knowledge about the means $\{\theta_{i}\}$), has to choose only one of these arms to play and accrue a reward at each time step, given the reward history so far observed. The target of the player is to maximize the rate of gaining reward as the time horizon $T$ grows. In many applications of online decision-making problems, such as Internet advertising and recommendation systems, the MAB model is widely assumed with numerous algorithms proposed for efficient reinforcement learning, e.g., Zhao [38], Lattimore and Szepesvári [22]. In a recommendation system, the vendor desires to know which item an online customer is looking for with the most interest. When we suggest a product and the customer makes a purchase, we receive a positive reward. Conversely, if the customer decides not to buy the recommended item, our reward may be either zero or negative, reflecting a potential waste of resources. Therefore the target of a recommendation system is to guess which items are most wanted by each customer, and the problem can be translated into the MAB model directly.
@@
To address the key issue in the conflict between exploration and exploitation, we note that it is impossible for the player to exactly know the expected values of the reward distributions within any finite horizon of time. The only information the player possesses is the (local) reward history, and the player has to learn from this to make a decision at each time step. Based on the limited historical information, the player has to choose between two options. The first option is to explore the arms with a lower average reward, as there is always a possibility that the best arm may not give the best average reward in the past sample path. The second option is to exploit the arm with the highest average reward offered so far. Robbins [28] considered a simple two-armed bandit problem and showed that there is a policy to achieve the best average reward asymptotically as time goes to infinity. Many years later, Lai and Robbins [20] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the benchmark case that the player already knows which arm has the highest reward mean from the very beginning. Given a set of $K$ arms with reward distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$ with means $\{\theta_{i}\}_{i=1}^{K}$ and a policy $\pi$, the regret ~$\tilde{\mathcal{R}}_{\pi}^{\mathcal{F}}(T)$ is defined as (without loss of generality, we can assume that $\theta_{1}\geq\theta_{i}$ for $1\leq i\leq K$)
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=\sum_{\theta_{i}<\theta_{1}}E[s_{i}(T)]\Delta_{i},$ (1)
@@
Strikingly, Lai and Robbins [20] proved an asymptotic lower bound on regret, with the growth rate of $O(\log T)$, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown. They also proposed an asymptotically optimal policy for a family of distributions, under which the best regret can be achieved asymptotically, both the logarithmic order and the leading coefficient. Later Burnetas and Katehakis [6] proved a generalized result under non-parametric settings. Suppose $\mathcal{H}$ is a set of distributions on $\mathbb{R}$. We call a bandit $\mathcal{F}$ (considered as a set of reward distributions of all arms) non-trivial, if there is at least one distribution (re-index it to $1$ if necessary) $F_{1}\in\mathcal{F}$ satisfying $\theta_{1}>\theta_{i}$ for all $i\in\{2,\dots,K\}$. They proved that for any policy $\pi$ s.t. for each fixed nontrivial $\mathcal{F}\subset\mathcal{H}$ (this policy condition is called consistency over $\mathcal{H}$)
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=o(T^{a}),\ \forall a>0,$ (3)
we have that
$\lim\inf\frac{\mathbb{E}[s_{i}(T)]}{\log T}\geq\sup\left\{\frac{1}{\text{KL}(F_{i},Y)}:Y\in\mathcal{H}\ \text{and}\ \mathbb{E}Y>\theta_{i}\right\},$ (4)
where $\text{KL}(\cdot,\cdot)$ denotes the Kullback-Leibler distance between two distributions. This theorem demonstrates that for any consistent strategy under a given family of distributions, the regret will not be lower than the logarithmic growth. Auer et al. [3] proposed a class of policies named and featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $[a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player. Their UCB1 and UCB1-Tuned polices perform very well for the classical MAB problem with a theoretically proven regret bound for any finite time horizon. The UCB policies are widely applied in machine learning development, e.g., the recommendations systems [32, 16], and in various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [19], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [15].
@@ 1.1 Heavy-tailed MAB Model
The class of heavy-tailed distributions is a concept in contrast to the light-tailed case. Light-tailed (or, sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation, while heavy-tailed distributions have a relatively high probability of producing a large deviation. There are several popular light-tailed probability distributions, such as Bernoulli, Gaussian, Laplacian, and Exponential. Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions [10]. Formally, a random variable $X$ is called light-tailed if there exists some $u_{0}>0$ such that its moment-generating function is well-defined on $[-u_{0},u_{0}]$ [10]:
@@
The above condition is equivalent to [10]
@@
where $M^{(2)}(\cdot)$ denotes the second derivative of $M(\cdot)$. Note that the above upper bound on $M(u)$ has a form of the moment generating function of the Gaussian distribution, thus the name “sub-Gaussian” has been adopted.
The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution might take large values with higher probabilities, thus its moment-generating function does not exist (even locally). Compared to the light-tailed distributions, the heavy-tailed are harder to learn in terms of the rank of their means. Bubeck et al. [5] proposed the robust UCB algorithm, under the assumption that there exists a known $p>1$ such that $\mathbb{E}[|X_{i}|^{p}]$ exists and is upper bounded by a known parameter, for all $1\leq i\leq K$. Robust UCB shows significant progress for the heavy-tailed MAB model by achieving the optimal logarithmic order of regret growth. Later on, Lattimore [21] developed a new UCB policy, which is referred to as the scale free algorithm. This algorithm removes the general assumption of a known upper bound on the $p$-th moment. Instead, it allows the variance and the squared fourth moment to scale freely as long as their ratio (kurtosis) is bounded by a known parameter. But this assumption restricts $p$ to be no less than $4$ (thus cannot be too heavy-tailed). In this paper, we extend Lattimore’s work to any $p>1$ with a tighter regret upper bound in the case of $p=4$ and low discrimination where arm means are close and learning the arm rank becomes challenging.
From this point on, we assume that the player knows a constant $C_{p,q}$ for some $p$ and $q$ ($p>q>1$) such that
@@
The above inequality is a direct generalization of the assumption on kurtosis in Lattimore [21] where $p=4$ and $q=2$. For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function [2]
@@
However, the moment under this distribution is a function of $\sigma$; when $\sigma$ increases, the moment of any fixed order also increases, if it exists. The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment [5]. Therefore, for an arm with an unknown $\sigma$, we cannot apply the robust UCB policy. Furthermore, the learning efficiency can be improved if a larger $p$ can be used (tighter bound on the tail, but the corresponding moment bound becomes harder to know or compute a priori), it is desired to eliminate the need for any prior knowledge regarding the specific values or bounds of the moments. This motivates Lattimore [21] to consider a scale free algorithm. After we generalize Lattimore’s work, we will give a correction to a minor flaw in a lemma of Bubeck et al. [5] which subsequently caused a small error in the computation of regret upper bound in Lattimore [21].
1.2 Related Works
@@
The theorems of the frequentist MAB have been developed for a long time and various milestones have been made since Lai and Robbins [20]. In contrast to the Bayesian model, the frequentist MAB does not assume a priori knowledge of the initial probability or state of the system but learns the core parameters solely through the past sampling history, e.g., computing an upper confidence bound as a function of the frequency that an arm is selected and the values of the observed samples. Such a way of decision-making in balancing exploration with exploitation contrasts to the Bayesian MAB where the tradeoff is usually addressed by dynamic programming for the time-varying characteristics of the system or the observation model [12]. After the UCB1 policy by Auer et al. in [3], many other policies of the UCB-type have been proposed in the literature. Maillard et al. [26] directed their attention to the prospect of KL divergence between probability distributions. They introduced the KL-UCB algorithm, which notably enhanced the regret bound. This improvement was achieved under the assumption of rewards being bounded and having a known upper bound. Furthermore, KL-UCB policy achieves the asymptotic lower bound of regret growth for this class of rewards given in Burnetas and Katehakis [6]. Kaufmann et al. [18] combined the Bayesian method with the UCB class and proposed Bayes-UCB for bounded rewards with a known bound, achieving a theoretical regret bound similar to that of KL-UCB. Numerical experiments demonstrated the high efficiency of learning by Bayes-UCB for Gaussian reward distributions. A survey of such UCB policies can be found in Burtini et al. [7]. These policies were neither implemented nor proved to achieve the logarithmic regret growth in the case of unbounded reward distributions. In 2011, the first author of this paper proposed UCB1-LT with a proven bound on the regret growth of the logarithmic order for the class of light-tailed reward distributions [24], filling the vacancy for the case of unbounded light-tailed reward distributions. Together with the extended robust UCB for the heavy-tailed class focused in this paper, we provide a relatively complete picture of UCB policies for general reward distributions. After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
@@
For the case of heavy-tailed reward distributions, we have mentioned that Bubeck et al. [5] proposed the robust UCB policy under the assumption that an upper bound of the moments is known. The robust UCB policy uses different mean estimators rather than the empirical one and remarkably achieves the logarithmic regret growth rate. Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator [8]. The truncated mean estimator requires the knowledge of an upper bound of origin moments, which is not very effective in the case that the reward mean is far away from zero. The median-of-means estimator considers the (central) moment and thus patches up the deficiency of the truncated mean estimator: the mean estimation will not be affected when the reward distributions are changed by translations. The Catoni mean estimator is more complicated and can only be used when variances of the reward distributions exist ($p\geq 2$). But when applied in the robust UCB policy, it gives a much smaller regret coefficient than the other two estimators. Recently Chen et al. [11] showed a result of applying Catoni mean estimator to the case that $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators. Finer estimations when extending the Catoni mean estimator to the case of $p<2$ are interesting for future investigations. Several recent papers also considered the heavy-tailed MAB model and made significant progresses. Wei and Srivastava [36] proposed the robust MOSS policy with logarithmic regret growth. Agrawal et al. [1] established a policy called $\text{KL}_{\inf}$-UCB, which achieves a logarithmic upper bound of regret very close to the asymptotic lower bound proved by Burnetas and Katehakis [6]. For other non-classical MAB models, the heavy-tailed model has also been considered [37, 30]. However, all these policies do not discard the knowledge requirement of an upper bound of moments, which has been removed from the extended robust UCB policy proposed in this paper.
@@
There are also policies not following the idea of UCB. The $\varepsilon$-greedy policy is well-studied, and many policies are proposed using its idea, e.g., the constant $\varepsilon$-decreasing policy [34], GreedyMix [9] and $\varepsilon_{n}$-greedy [3]. The Deterministic Sequencing of Exploration and Exploitation (DSEE) policy [25, 33] achieves the logarithmic regret for both cases of single player and multiple players (with the knowledge on moment upper bounds). Other recent progresses on variants of MAB, such as adversarial bandits, contextual bandits and linear bandits, can be found in Lattimore and Szepesvári [22, 13, 17].
2 The Heavy-Tailed Distribution

### v4

@@ Abstract
The multi-armed bandit (MAB) problem is a widely studied model in the field of operations research for sequential decision making and reinforcement learning. This paper mainly considers the classical MAB model with the heavy-tailed reward distributions. We introduce the extended robust UCB policy, which is an extension of the pioneering UCB policies proposed by Bubeck et al. [5] and Lattimore [22]. The previous UCB policies require some strict conditions on the reward distributions, which can be hard to guarantee in practical scenarios. Our extended robust UCB generalizes Lattimore’s seminary work (for moments of orders $p=4$ and $q=2$) to arbitrarily chosen $p>q>1$ as long as the two moments have a known controlled relationship, while still achieving the optimal regret growth order $O(\log T)$, thus providing a broadened application area of the UCB policies for the heavy-tailed reward distributions. Furthermore, we achieve a near-optimal regret order without any knowledge of the reward distributions as long as their $p$-th moments exist for some $p>1$.
@@ 1 Introduction
The multi-armed bandit problem is an important topic in operations research where a player aims to identify and select the best arm to operate such that the long-term expected reward is maximized. The player’s action at each time involves balancing the trade-off between exploration and exploitation, i.e., choosing an arm to learn its statistics vs. choosing an arm appearing to be best based on past observations. Under resource constraints, the way we acquire knowledge becomes crucial in maximizing the long-term reward rate in modern AI systems [31]. Since knowledge can generally be treated as data, the struggle between exploration (sampling) and exploitation (consuming) as the prototypical dilemma in reinforcement learning is expressly reflected in an MAB model [28, 21, 3, 24, 5]. Specifically, the original stochastic model of the multi-armed bandit problem considers a bandit machine with several arms; each arm $i$ offers a random reward according to an unknown distribution $X_{i}$ when pulled by the player (or agent). Each reward distribution has a mean, say $\theta_{i}$. The player who has limited knowledge about the underlying reward distributions (no knowledge about the means $\{\theta_{i}\}$), has to choose only one of these arms to play and accrue a reward at each time step, given the reward history so far observed. The aim of the player is to maximize the reward rate that accrues as the time horizon $T$ grows. In many applications of online decision-making problems, such as Internet advertising and recommendation systems, the MAB model is widely assumed, with numerous algorithms proposed for efficient reinforcement learning, e.g. Zhao [38], Lattimore and Szepesvári [23].
@@
To address the key issue in the conflict between exploration and exploitation, we note that it is impossible for the player to exactly know the values of reward means within any finite horizon of time. The only information the player possesses is the (local) reward history, and the player has to learn from this to make a decision at each time step. Based on the limited historical information, the player has to choose between two options. The first option is to explore a suboptimal arm with a lower average reward, as there is always a possibility that the best arm may not give the best average reward in the past sample path. The second option is to exploit the arm with the highest average reward offered so far. Robbins [28] considered a simple two-armed bandit problem and showed that there is a policy to achieve the best average reward asymptotically as time goes to infinity. Many years later,Lai and Robbins [21] proposed a much stronger performance measure referred to as regret, which represents the expected total loss compared to the benchmark scenario in which the player already knows which arm has the highest reward mean from the very beginning. Given a set of $K$ arms with reward distributions =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{K}\}$ with means $\{\theta_{i}\}_{i=1}^{K}$ and a policy $\pi$, the regret ~$\tilde{\mathcal{R}}_{\pi}^{\mathcal{F}}(T)$ is defined as (without loss of generality, we can assume that $\theta_{1}\geq\theta_{i}$ for $1\leq i\leq K$)
$\mathcal{R}_{\pi}^{\mathcal{F}}(T)=\sum_{\theta_{i}<\theta_{1}}\mathbb{E}[s_{i}(T)]\Delta_{i},$ (1)
@@
Strikingly, Lai and Robbins [21] proved an asymptotic $O(\log T)$ lower bound on the regret growth rate, under the assumption that the probability density/mass function of each reward distribution has the form $f(\cdot;\theta)$, where $f$ is known but parameter $\theta$ is unknown. They also proposed an asymptotically optimal policy for a family of distributions, under which the best regret can be achieved asymptotically, both the logarithmic order and the leading coefficient. Auer et al. [3] proposed a class of policies featured by UCB (upper confidence bound), achieving logarithmic regret if the reward distributions have finite support on a closed interval $[a,b]\subset\mathbb{R}$ with $a,b\in\mathbb{R}$ known to the player. Their UCB1 and UCB1-Tuned polices perform very well for the classical MAB problem with a theoretically proven regret bound for any finite time horizon. UCB policies are widely applied in machine learning development (e.g. recommendation systems [32, 17]) and various reinforcement learning algorithms such as Monte Carlo Tree Search (MCTS) [20], which is frequently used in gaming AIs [27] and forms a key step in the design of AlphaGo Zero [16].
@@ 1.1 Heavy-Tailed MAB Model
The class of heavy-tailed distributions is a concept in contrast to the light-tailed case. Light-tailed (or sub-Gaussian) distributions usually have only a relatively small probability of producing a significantly shifted random sample from its expectation, while heavy-tailed distributions have a relatively high probability of producing a large deviation. There are several popular light-tailed probability distributions, such as Bernoulli, Gaussian, Laplacian, and Exponential. Specifically, the class of light-tailed distributions requires the (local) existence of the moment-generating function of the associated random variable and is therefore referred to as the locally sub-Gaussian distributions Chareka [10]. Formally, a random variable $X$ is called light-tailed if there exists some $u_{0}>0$ such that its moment-generating function is well-defined on $[-u_{0},u_{0}]$ [10]:
@@
The above condition is equivalent to [10]:
@@
where $M^{(2)}(\cdot)$ denotes the second derivative of $M(\cdot)$. Note that the above upper bound on $M(u)$ has a form of the moment-generating function of the Gaussian distribution, thus the name “sub-Gaussian” has been adopted.
The terminology “heavy-tailed”, opposite to “light-tailed”, implies that the reward distribution can produce large values with high probabilities, leading to the nonexistence of its moment-generating function (even locally). Compared to the light-tailed class, the heavy-tailed one is harder to learn in terms of the rank of their means. Bubeck et al. [5] proposed the robust UCB algorithm, under the assumption that there exists a known $p>1$ such that $\mathbb{E}[|X_{i}|^{p}]$ exists and is upper bounded by a known parameter, for all $1\leq i\leq K$. Robust UCB shows significant progress for the heavy-tailed MAB model by achieving the optimal logarithmic order of regret growth. Subsequently, Lattimore [22] developed a new UCB policy referred to as the scale free algorithm. This algorithm removes the general assumption of a known upper bound on the $p$-th moment. Instead, it allows the variance and the squared fourth moment to scale freely as long as their ratio (kurtosis) is bounded by a known parameter. But this assumption restricts $p$ to be no less than $4$ (not too heavy-tailed). In this paper, we extend Lattimore’s work to any $p>1$ with a tighter regret upper bound in the case of $p=4$ with low discrimination where arm means are close and learning the arm rank becomes challenging.
First, we assume that the player knows a constant $C_{p,q}$ for some $p$ and $q$ ($p>q>1$) such that [^1] By Jensen’s inequality, the existence of the $p$-th moment implies the existence of the $q$-th moment for any $1<q<p$.
@@
The requirement of knowledge of $C_{p,q}$ can be removed with an arbitrarily small sacrifice of the regret order as seen in Sec. 2.1. The above inequality is a direct generalization of the assumption of kurtosis in Lattimore [22] where $p=4$ and $q=2$. For a specific example, suppose that $X$ has a Pareto distribution of type III with cumulative distribution function Arnold [2]
@@
However, the moment of this distribution is a function of $\sigma$. When $\sigma$ increases, the moment of any fixed order also increases, if it exists. The robust UCB policy requires the knowledge of $p$ and an upper bound of the $p$-th moment Bubeck et al. [5]. Therefore, for an arm with an unknown $\sigma$, we cannot apply the robust UCB policy. Furthermore, learning efficiency can be improved if a larger $p$ can be used (tighter bound on the tail, but the corresponding moment bound becomes harder to know or compute a priori). It is thus better to eliminate the need for any prior knowledge regarding the specific values or bounds on the moments. This motivates Lattimore [22] to consider a scale free algorithm. After we generalize Lattimore’s work, we will give a correction to a minor flaw of a lemma in Bubeck et al. [5] that subsequently caused a small error in the computation of the regret upper bound in Lattimore [22].
1.2 Related Work
@@
The theorems of the frequentist MAB have been developed for a long period of time and various milestones have been made since Lai and Robbins [21]. In contrast to the Bayesian model, the frequentist MAB does not assume a priori knowledge of the initial probability or state of the system but learns the core parameters solely through the past sampling history, e.g., computing an upper confidence bound as a function of the frequency that an arm is selected and the values of the observed samples. This way of decision-making in balancing exploration with exploitation contrasts with the Bayesian model in which the tradeoff is usually addressed by dynamic programming for the transitional characteristics of the system or the observation model [13]. After the UCB1 policy by Auer et al. [3], many other policies of the UCB class have been proposed in the literature. Maillard et al. [26] directed their attention to the prospect of KL divergence between probability distributions. They introduced the KL-UCB algorithm, which notably tightened the regret bound. This improvement was achieved under the assumption of bounded rewards with a known bound. Furthermore, KL-UCB policy achieves the asymptotic lower bound of regret growth for this class of rewards given in Burnetas and Katehakis [6]. Kaufmann et al. [19] combined the Bayesian method with the UCB class and proposed Bayes-UCB for bounded rewards with a known bound, achieving a theoretical regret bound similar to that of KL-UCB. Numerical experiments demonstrated the high efficiency of learning by Bayes-UCB for Gaussian reward distributions. A survey of such UCB policies can be found in Burtini et al. [7]. These policies were neither implemented nor proved to achieve the logarithmic regret growth in the case of unbounded reward distributions. In 2011, the first author of this paper proposed UCB1-LT with a proven bound on the regret growth of the logarithmic order for the class of light-tailed reward distributions Liu and Zhao [24], filling the gap for the case of unbounded light-tailed reward distributions. Together with the extended robust UCB for the heavy-tailed class focused in this paper, we provide a relatively complete picture of UCB policies for general reward distributions. After 2011, Bubeck and Cesa-Bianchi [4] extended UCB1-LT to $(\alpha,\psi)$-UCB under a more general assumption that the reward moment-generating function is bounded by a convex function $\psi$ for all $u\in\mathbb{R}$.
@@
For the case of heavy-tailed reward distributions, we have mentioned that Bubeck et al. [5] proposed the robust UCB policy under the assumption that an upper bound of the moments is known. The robust UCB policy uses different mean estimators rather than the empirical one and remarkably achieves the logarithmic regret growth rate. Under certain regularity conditions, the robust UCB policy may adopt three different mean estimators: the truncated mean estimator, the median-of-means estimator, and the Catoni mean estimator Catoni [8]. The truncated mean estimator requires the knowledge of an upper bound of origin moments, which is not very effective in the case that the reward mean is far away from zero. The median-of-means estimator considers the (central) moment and thus patches up the deficiency of the truncated mean estimator: the mean estimation will not be affected when the reward distributions are transformed by translations. The Catoni mean estimator is more complicated and can only be used when the reward distributions have variances ($p\geq 2$). But when applied in the robust UCB policy, it gives a much smaller regret coefficient than the other two estimators. Recently, Chen et al. [11] showed a result of applying Catoni mean estimator to the case of $p<2$, but the performance was not proven to be better than the truncated mean and the median-of-means estimators. Finer estimations when extending the Catoni mean estimator to the case of $p<2$ are interesting for future investigations. Several recent papers also considered the heavy-tailed MAB model and made significant progresses. Wei and Srivastava [36] proposed the robust MOSS policy with logarithmic regret growth. Agrawal et al. [1] established a policy called $\text{KL}_{\inf}$-UCB, which achieves a logarithmic upper bound on regret very close to the asymptotic lower bound proved by  Burnetas and Katehakis [6]. For other non-classical MAB models, the heavy-tailed model has also been considered [37, 30]. However, these policies do not eliminate the knowledge requirement of an upper bound of moments, which has been removed from the extended robust UCB policy proposed in this paper. Note that Genalti et al. [12] considered the case where no knowledge of the moments is available except for the existence of moments of order greater than $1$, but a relatively strong assumption of the distribution (a larger probability mass on the negative semi-axis) was needed to achieve logarithmic regret.
@@
There are also policies not following the idea of UCB. The $\epsilon$-greedy policy was well studied, and many policies were proposed using its idea, e.g., the constant $\epsilon$-decreasing policy [34], GreedyMix [9], and $\epsilon_{n}$-greedy [3]. The Deterministic Sequencing of Exploration and Exploitation (DSEE) policy [25, 33] also achieves logarithmic regret for both cases of single player and multiple players (with knowledge of upper bound on moments). Other recent progresses on variants of MAB, such as adversarial bandits, contextual bandits, and linear bandits, can be found in [23, 14, 18].
2 Heavy-Tailed Reward Distributions